# Emergence of Quantised Representations Isolated to Anisotropic Functions

### v2

@@ Abstract
This paper presents a novel methodology for determining representational alignment, which builds upon the existing Spotlight Resonance method. Particularly, this new tool is used to gain insight into how discrete representations can emerge and organise in autoencoder models, through a controlled ablation study in which only the activation function is altered. Using this technique, the validity of whether function-driven symmetries can act as implicit inductive biases on representations is determined. Representations are found to tend to discretise when the activation functions are defined through a discrete algebraic permutation-equivariant symmetry. In contrast, they remain continuous under a continuous algebraic orthogonal-equivariant definition. This confirms the hypothesis: algebraic symmetries of network primitives can carry unintended inductive biases which produce task-independent artefactual structures in representations. The discrete symmetry of contemporary forms is shown to be a strong predictor for the induction of discrete representations transformed from otherwise continuous structures — a quantisation effect. This motivates further reassessment of functional forms in common usage. Moreover, this supports a general causal model for one mode in which discrete representations may form, and could constitute a prerequisite for downstream interpretability phenomena, including grandmother neurons, discrete coding schemes, general linear features and possibly Superposition. Hence, this tool and proposed mechanism for the influence of functional form on representations may provide insights into emergent interpretability research. Finally, preliminary results indicate that quantisation of representations appears to correlate with a measurable increase in reconstruction error, reinforcing previous conjectures that this collapse can be detrimental.
@@ 1 Introduction
1.1 How Functional Form Choices May Result in Representational Inductive Biases
Many of the foundational functions in contemporary deep learning produce maps that differ as the direction changes; a property termed as anisotropy. This angular variation of the functions is hypothesised to make specific directions distinct, which the network then alters its representations about through optimisation. Hence, an angularly varying map could be expected to induce shaped angular representations. It is this connection between functional forms and representations which shall be explored in this work.
@@ 1.1 How Functional Form Choices May Result in Representational Inductive Biases
In particular, contemporary functional forms tend to make distinct the standard basis’ vectors, $\hat{e}_{i}$, and in such a way that each basis vector’s distinction is equal. This particular form of anisotropy can be categorised through their shared, underlying permutation symmetry construction Godfrey et al. (2023); Bird (2025a). This can now act as a defining characteristic relation for most contemporary functional forms. This is often due to their elementwise application, but appears more generally. The specific relation depends on the primitive in question; for most activation functions, this is an algebraic permutation equivariance Godfrey et al. (2023); Bird (2025a), as displayed in Eqn. 1. The hypothesis is deepened by suggesting that the maximal symmetries to which functional forms belong are a strong predictor of representational structure.
@@
The Spotlight Resonance Method Bird (2025b) has already empirically demonstrated how such activation functions can produce a tendency for representational alignments about the distinguished directions. This was achieved through transforms, and altering the number of the functional form’s distinguished directions, with corresponding changes in the representations mirroring the transformed structure of the distinguished directions being observed. The conclusion was that functional forms stimulate this tendency towards representational alignment, producing emergent task-agnostic alignment structure in representations, and hence establishing a causal relationship between the two. The observed structure consisted of a tendency for a greater density of representations to cluster about the functional form’s distinguished basis. This work also indicated the presence of ’grandmother neurons’, indicating how many clusters represent a distinct semantic and also associated their alignment with functional form choices. However, it did not establish whether discretisation itself was a phenomenon of this choice of form.
1.2 Motivations for This Study
@@
Isotropic Deep Learning Bird (2025a) takes the consequences of such observations a step further. The Spotlight Resonance Method determined that task-agnostic structure in representations is associated with a functional form’s distinguished directions. The Isotropic Deep Learning approach proposes that since this structure transforms with the functional form, the anisotropic functional form choices may be the fundamental causal mode that induces the array of interpretable structures to begin with. In effect, anisotropic primitives provide an absolute frame, for each representation space, about which the representations organise and discretise into distinct clusters. Hence, questioning whether the ubiquity of anisotropy in contemporary deep learning constitutes an underlying origin for many such interpretable and emergent-structure phenomena, which are commonly observed — and whether this is beneficial or not. Furthermore, determining the validity of such a hypothesis would indicate whether these phenomena are truly fundamental to deep learning or stimulated by, and contingent on, the specific choices in functional form. This could have significant implications for the approach of network interpretability. Confirmation of this hypothesis would also demonstrate that such choices do carry inherent and unappreciated biases with their implementation.
@@
This position motivates the comparative study of this work. Determining whether one can control and demonstrate the emergence and elimination of structure through broadening deep learning’s primitives to other form-defining choices, each providing its own inductive bias. One such proposed choice Bird (2025a) is by redefining primitives through an orthogonal algebraic-equivariance, displayed in Eqn. 2. This is termed an ‘isotropic’ definition of forms, and is a supergroup-symmetry of the original permutation symmetry, $S_{n}\subset\mathrm{O}\left(n\right)$. This defining choice of isotropic primitives is interesting, since it can eliminate all form-distinguished directions in a network. Therefore, functions developed to adhere to this modified symmetry definition can be used as a comparative mode to determine whether functional forms can induce these interpretable structures and indicate whether downstream alignment phenomena are attributable to anisotropic choices. If this is the case, it would demonstrate that these differing forms do carry unintended inductive biases, which would warrant major reconsideration of the forms used in standard practice.
@@
The hypothesis is that since isotropic functional forms are devoid of distinguished directions, they may consequently eliminate any tendency towards task-agnostic representational collapse aligned along these directions. This was one prediction made by the author regarding why isotropic primitives are speculated to be beneficial and is discussed further below. This motivates the importance of the present study as a starting point for probing this hypothesised induced representational collapse and its consequences.
In particular, anisotropic forms may lead to a tendency for representations to overly cluster around distinguished directions. This is due to the anisotropic functional form’s nature, where some angular arrangements may become overrepresented through optimisation due to their irregularity. This task-agnostic over-density bias may be detrimental, as representations may tend to collapse towards single directions, such as aligned or anti-aligned, and consequently lose informative representational degrees of freedom, potentially damaging a network’s internal expressiveness of otherwise continuous structure. Hence, anisotropy may result in a quantising effect on representations, and thus also the semantics being represented — resulting in a limited representational capacity and neural coding scheme. Consequently, isotropic primitives are expected to remove this form-induced task-agnostic discretising bias on representations.
This may occur through several modes, hypothesised to stem from the underlying symmetry. To begin with, these symmetries are able to interact with the network structure and therefore its dynamics. This differs from symmetries in end-to-end networks, where their representations are designed to transform with the symmetry, resulting in the entire model being either equivariant or invariant to its effect. This is not the case in general networks, where a symmetry action does not pass cleanly through or identically cancel with layers. In general networks, a layer may ‘absorb’ a symmetry action in a non-invariant/equivariant manner, or have an interaction with its differing symmetry, meaning the symmetry can, in principle, ‘act’ on the network and hence produce a realisable effect. This is argued to result in various inductive biases that are considered underappreciated in significance. For example, the permutation-like symmetries about the standard basis result in an orthant partitioning of the space. Additionally, depending on the particular permutation symmetry, sets of these orthants may be analytically degenerate under hyperoctahedral transforms. The network’s optimisation will then shape representations about these orthants of differing and equal mapping, to achieve the desired computation. This results in representations that organise differently due to maps defined by the symmetry of the form, and in their extreme, may cluster over particular regions, producing a quantisation effect on representations. This is considered a general mode through which function-driven symmetries may influence representations as optimisation responds to their structure. The general approach is to study how various function-driven symmetries interact and act on the network and its representations, determining the various biases induced.
This mechanism would aid in explaining why the SRM study Bird (2025b) detected axis-aligned structure, since incidental alignment with directions people choose to define and observe as ‘individual neurons’ would have a vanishingly small likelihood of occurring in high-dimensional spaces. This is because there are vastly more orientations that do not align with the standard basis than those that do in high-dimensional spaces. However, if the analytic properties of permutation-based functions produce maps which differ along these directions, then the symmetry breaking of the space would single out such directions and make substantial alignment less peculiar. Therefore, it indicates an unintended bias of the model. A comparison between anisotropic and isotropic forms can directly probe such questions by adding and removing any distinction between directions.
Quantisation is argued to have potential pathological effects on the network. One heuristic is that the underlying distributions might be better represented continuously, as this aligns with the frequently continuous real-world semantics being represented, such as broad object morphologies and arangments, multivariate signals, lighting and shades, continuous expressions, modelling distributions, etc [^1] This aspect independently converges on similar motivation to Sabour et al. (2017)’s vector-valued routed capsule models but significantly diverges in terms of lack of capsules and routing and it is foremost concerning general architectures (potentially under reselection). Substantial differences arise also in the reconsideration of, and interactions between, sets of primitices, additionally the generality of the scope of such functional forms outside of just normalisation, alongside the investigation and leveraging of implicit inductive biases regarding various symmetries constructions. Therefore, this independent approach constitutes a considerably broader consideration of the impacts and generality of such forms and their symmetry-defining approach in deep learning Bird (2025a).. Therefore, enabling latent variables and representations to be continuous and potentially more interpolatable at various levels through a network, instead of an imposed task-agnostic structure, may be beneficial for many tasks. This may increase representational capacity and organisation of the representations.
This connects to the coding scheme of the network. Axis-aligned representations, previously observed to frequently emerge Bird (2025b), approximate a local coding scheme for the network. This is where one neuron corresponds to one semantic. This coding scheme has decreased concept-interference, but has a much lower representational capacity than other coding schemes. Models have been demonstrated to produce other interference-capacity balances with various input sparsities Elhage et al. (2022); Bird (2025b), suggesting that axis-alignment is not an appropriate universal inductive bias — further motivating the study of alternative functional form inductive biases and their effects on representations.
A functional form that does not distinguish particular directions may therefore not deter such continuous and interpolatable structures present in the dataset, and their preservation may be advantageous to the network. Yet, sometimes discretisation may be appropriate, such as in classification networks, where abstracting away redundant detail can be beneficial — though in autoencoding models this would not be an expectation. In any case, rigidly imposing this through a task-agnostic, function-driven, inductive bias about a specific arrangement might be detrimental. For example, when reducing the dimensionality of a dataset inherently containing discrete representations via a linear layer, to some arbitrary lower dimension, one would not expect it always appropriate for those discrete directions to be best represented through an alignment to an orthonormal basis or other imposed geometry for that space — as these are as arbitrary to the bottleneck chosen.
Moreover, it is essential to note that isotropic networks still enable such clustering. They retain the ability through their bias terms, but they are suggested not to disproportionately encourage it through task-agnostic biases; instead, it would likely reflect a task-necessitated collapse. Hence, there is no reason to expect isotropy in representational distributions, as the underlying data or latent layer may include discretised aspects that arise independently of model effects. Therefore, this is not a test of whether distributions are anisotropic or isotropic; instead, it is the investigation of whether the emergent structure specifically about a distinguished direction can be turned on or off through functional form choices. Therefore, isotropic networks may be expected to produce emergent structures that are more representative of the underlying data, rather than being dependent on functional forms. Hence, proposing that these choices carry representational inductive biases, where isotropy’s frame independence may be a less arbitrary, and thus more preferable and natural choice.
Other speculative pathological modes are attributed to the hypothesised functional form quantisation. For example, discrete representations may be particularly susceptible to perturbations which move activations outside of the cluster, into an in-between region of representation space. Maps from these unpopulated regions may be poorly trained and fragile, due to their training-time rarity, and therefore highly unpredictable in their map. This may suggest that such a quantising inductive bias decreases adversarial robustness, whereas removing this bias through isotropy may be advantageous. However, this speculation is not explored within this work.
This paper aims to assess the validity of the former prediction. The motivating questions are: do discrete-symmetry defined functional forms tend to produce approximately discrete representations, quantising an otherwise underlying continuous data structure, whilst continuous-symmetry defined functional forms may tend to preserve a more continuous representation. Furthermore, is the prior construction potentially pathological in nature? Do continuous representations tend to improve performance, since quantised representations may harm expressiveness? Therefore, determining the validity of this prediction may provide crucial insight into the potential of Isotropic deep learning and the emergence of downstream interpretability structure, namely alignment, grandmother neurons, general interpretable directions, discrete codes, and possibly Superposition Elhage et al. (2022) and Neural Collapse Papyan et al. (2020).
Consequently, this study can be considered to be determining a more foundational prerequisite phenomenon than the Spotlight Resonance Method observed and causally explained. This prior method elucidated how functional forms can causally influence alignment structure. In contrast, this work investigates whether forms can lead to the emergence of any induced task-agnostic structure. Hence, determining whether current forms induce representational quantisation may indicate a causal mechanism for all downstream phenomena. This study supports the hypothesis that such choices may be significant representational inductive biases and worth consideration.
Most importantly, this study aims to clarify the link between functional form algebraic symmetries and their impact on representations. It will first be demonstrated that symmetry is a key predictor of the emergent task-agnostic structure in representations. This encourages an exploration into the effects of various symmetry groups and possibly a reconsideration of functional form primitives, network-wide, if alternative symmetries are beneficial. This is because symmetry can act as a classifying and definitional approach to functional forms.
Additionally, through this approach, some preliminary evidence is provided that such representational collapses may be pathological in nature, increasing task error. However, this study is not intended to elucidate the exact mode of the pathology, such as those proposed above, so the results remain preliminary in this regard. Primarily, this study investigates how representations may quantise due to functional form, rather than studying how they may align due to functional form.
@@ 1.3 Methodological Approach
The Spotlight Resonance Method hypothesised a hierarchy of anisotropic influences, particularly through the triad discussed earlier. These interacting distinguished bases may be highly non-trivial to predict the representational alignment of. Therefore, the causal mechanism was established in a clean, isolated and minimalistic way through ablation studies of only activation functions to minimise conflating or obscuring influences. Establishing this base case of a causal mechanism from primitives to representations can then be expected to carry downstream through more general networks built on these primitive influences. The author’s approach of establishing such a causal mechanism through a minimalistic approach, which can then be expected to apply more generally, will be repeated in our methodology.
@@
Instead of rotating the privileged basis or altering the number of distinguished directions to determine alignment of representations, this study will compare representations between the presence and absence of any distinguished directions to determine if this is a causal mechanism for the production of approximately discrete representations. Due to the expected systematic nature of this inductive bias, care was taken to prevent the unintentional reintroduction of anisotropy. To achieve this, distinguished directions will be removed network-wide to avoid these confounding and obscuring interactions. This is achieved by utilising isotropic primitives and an autoencoder model for reconstruction  [^2] The use of an autoencoder model prevents anisotropy through choices in the output layer representation. For example, one-hot encoding in classification outputs is also a choice in form defined through permutation symmetry; therefore, it could induce further spurious structure. A reconstruction task using autoencoders may therefore mitigate this and enable more data-representative and ‘natural’ embeddings. Moreover, it has been speculated Bird (2025b, a) whether the neural-collapse classification phenomenon Papyan et al. (2020) is related to the proposed mechanisms. So autoencoder reconstruction provided a suitable minimalistic setting, free from these influences to study form-induced representational quantisation., except for an activation function which can selectively reintroduce anisotropy through changes to its form. Hence, this isolates the causality of the phenomenon to the activation functions.
@@
This requires two activation functions, which are highly similar in action but differ in the presence or absence of anisotropy. The isotropic position paper Bird (2025a) provides such a preliminary isotropic activation function, which enables this comparative analysis. An analogue of standard elementwise (anisotropic) tanh, displayed in Eqn. 3, can be compared with its orthogonal analogue Isotropic-Tanh, displayed in Eqn. 4 — each is displayed in multivariate form for comparison $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. These functions produce equivalent non-linear maps along the standard basis directions: $f_{\text{iso}}\left(\alpha\hat{e}_{i}\right)=f_{\text{aniso}}\left(\alpha\hat{e}_{i}\right)$ forall $\hat{e}_{i}$ and $\alpha$. Standard-Tanh is algebraically equivariant to the hyperoctahedral group (standard-basis permutation with sign-flips), $B_{n}$, whilst isotropic-tanh is algebraically equivariant to the orthogonal group, $\mathrm{O}\left(n\right)$, with $B_{n}\subset\mathrm{O}\left(n\right)$.
@@
This approach is also extended to a Leaky-ReLU Maas et al. (2013) family, which are modified to adhere to the following symmetries $S_{n}\subset\mathrm{O}\left(n\right)$. This is discussed further in App. B.4 and indicates that algebraic symmetries are the primary predictor, since all other algebraic properties of Leaky-ReLU substantially differ from the tanh-like family of Eqns. 3 and 4.
@@
Additionally, the PPP method does not produce any angular-falloff smearing effect related to the cone width of SRM. This modified method therefore enables higher-fidelity angular maps of representations, which may be beneficial in many circumstances. The following section provides a detailed overview of this tool.

### v3

@@ Abstract
This paper presents a novel methodology for determining representational structure, which builds upon the existing Spotlight Resonance method. This new tool is used to gain insight into how discrete representations can emerge and organise in autoencoder models, through a controlled ablation study in which only the activation function is altered. Using this technique, the validity of whether function-driven symmetries can act as implicit inductive biases on representations is determined. Representations are found to tend to discretise when the activation functions are defined through a discrete algebraic permutation-equivariant symmetry. In contrast, they remain continuous under a continuous algebraic orthogonal-equivariant definition. This confirms the hypothesis that the symmetries of network primitives can carry unintended inductive biases, which produce task-independent artefactual structures in representations. The discrete symmetry of contemporary forms is shown to be a strong predictor for the production of discrete representations emerging from otherwise continuous distributions — a quantisation effect. This motivates further reassessment of functional forms in common usage due to such unintended consequences. Moreover, this supports a general causal model for one mode in which discrete representations may form, and could constitute a prerequisite for downstream interpretability phenomena, including grandmother neurons, discrete coding schemes, general linear features and possibly Superposition. Hence, this tool and proposed mechanism for the influence of functional form on representations may provide insights into interpretability research. Finally, preliminary results indicate that quantisation of representations appears to correlate with a measurable increase in reconstruction error, reinforcing previous conjectures that this collapse can be detrimental.
@@ 1 Introduction
This introduction will begin by outlining how many of deep learning’s current functions are expected to influence the representations. This is hypothesised to constitute a function-driven emergent bias, motivating a comparative analysis against analogous functions adhering to differing symmetries. Finally, an approach is discussed for isolating such influences through a controlled ablation study.
1.1 Overview for How Functional Form Choices May Result in Biases
Many of the foundational functions in contemporary deep learning produce maps that differ as the direction changes; a property termed ‘anisotropy’. This angular variation of the functions is hypothesised to make specific directions distinct, which the network may then alter its representations about through optimisation. Hence, an angularly varying map may be expected to induce shaped angular representations. It is this connection between functional forms and representations which shall be explored in this work.
@@ 1.1 Overview for How Functional Form Choices May Result in Biases
In particular, contemporary functional forms tend to make distinct the standard basis’ vectors, $\hat{e}_{i}$, and in such a way that each basis vector’s distinction is equal. This particular form of anisotropy can also be denoted through a permutation symmetry Godfrey et al. (2023). This can be promoted to a defining characteristic relation underlying contemporary functional forms and categorised accordingly Bird (2025a). This permutation symmetry is often due to their elementwise application, but can appear more generally. The specific relation depends on the primitive in question; for most activation functions, this is an algebraic permutation equivariance. This permutation equivariance is displayed in Eqn. 1111This is assuming a map $\rho:\mathcal{G}\rightarrow\mathrm{GL}_{n}\left(\mathbb{R}\right)$ representation of the group, in this case one which permutes the standard basis. This form of representation will be assumed and implicit in all groups represented as matrices moving forward.. The hypothesis is that the maximal symmetries to which functional forms belong are a strong predictor of emergent representational structure.
@@
The Spotlight Resonance Method Bird (2025b) has already empirically demonstrated how such activation functions can produce a tendency for representational alignments about the distinguished directions. This was achieved through transformations to the standard basis, and also altering the number of distinguished directions through the functional form. It was found that corresponding changes in the representations occurred, which followed the transformed structure of the distinguished directions. The conclusion was that functional forms stimulate this tendency towards representational alignment, producing emergent task-agnostic alignment structure in representations, and hence establishing a causal relationship between the two. The observed structure showed a tendency for a greater density of representations to cluster around the functional form’s distinguished directions. This work also indicated the presence of ’grandmother neurons’, demonstrating how many clusters represent a distinct semantic. However, it did not establish whether discretisation itself was a phenomenon of this choice of functional form.
1.2 Motivating The Approach for this Study
@@
Isotropic Deep Learning Bird (2025a) takes the consequences of such observations a step further. The Spotlight Resonance Method determined that task-agnostic structure in representations is associated with a functional form’s distinguished directions. The Isotropic Deep Learning approach proposes that since this structure transforms with the functional form, the anisotropic functional form choices may be the fundamental causal mode that induces the array of interpretable structures to begin with.
In effect, anisotropic primitives provide an absolute frame, for each representation space, about which the representations organise and discretise into distinct clusters. Hence, it raises the question of whether the ubiquity of anisotropy in contemporary deep learning constitutes an underlying origin for many such interpretable and emergent-structure phenomena that are commonly observed. If this is the case, then it must be assessed whether this is beneficial or not. Furthermore, determining the validity of such a hypothesis would indicate whether these phenomena are truly fundamental to deep learning or stimulated by, and contingent on, the specific choices in functional form currently in use. This could have significant implications for the approach of network interpretability. Confirmation of this hypothesis would also demonstrate that such choices do carry inherent and unappreciated biases with their implementation.
@@
This position motivates the comparative study of this work. Determining whether one can control and demonstrate the emergence and elimination of structure through broadening deep learning’s primitives to other forms, particularly those defined through differing maximal symmetries. The hypothesis is that each may provide its own inductive bias. One such proposed choice is by redefining primitives through an orthogonal algebraic-equivariance Bird (2025a), displayed in Eqn. 2. This is termed an ‘isotropic’ definition for activation function forms, and is a supergroup-symmetry of the original permutation symmetry, $S_{n}\subset\mathrm{O}\left(n\right)$.
@@
Choosing this definition for primitives is interesting, since it can eliminate all form-distinguished directions in a network. Therefore, functions developed to adhere to this modified symmetry definition can be used as a comparative mode to determine whether functional forms can induce these interpretable structures and indicate whether downstream alignment phenomena are attributable to anisotropic choices. If this is the case, it would demonstrate that these differing forms do carry unintended inductive biases, which would warrant major reevaluation of the forms used in standard practice.
@@ 1.3 Methodological Approach
The Spotlight Resonance Method (SRM) Bird (2025b) hypothesised a hierarchy of anisotropic influences, particularly through the triad discussed earlier. These interacting distinguished bases may be highly non-trivial to predict the representational alignment of. Therefore, the causal mechanism was established in a clean, isolated and minimalistic way through ablation studies of only activation functions to minimise conflating or obscuring influences. Establishing this base case of a causal mechanism, from primitives to representations, can then be expected to carry downstream through more general networks built on these primitive influences. The SRM author’s approach of establishing such a causal mechanism in a minimalistic setting, which can then be expected to apply more generally, will be repeated in this methodology.
Instead of rotating the privileged basis or altering the number of distinguished directions to determine alignment of representations, this study will compare representations between the presence and absence of any distinguished directions to determine if this is a causal mechanism for the production of approximately discrete representations.
@@
Due to the expected systematic nature of this inductive bias, care was taken to prevent the unintentional reintroduction of anisotropy. To achieve this, distinguished directions will be removed network-wide to avoid these confounding and obscuring interactions. This is achieved by utilising isotropic primitives and an autoencoder model for reconstruction  [^2] The use of an autoencoder model prevents anisotropy through choices in the output layer representation. For example, one-hot encoding in classification outputs is also a choice in form defined through permutation symmetry; therefore, it could induce further spurious structure. A reconstruction task using autoencoders may therefore mitigate this and enable more data-representative and ‘natural’ embeddings. Moreover, it has been speculated Bird (2025b, a) whether the neural-collapse classification phenomenon Papyan et al. (2020) is related to the proposed mechanisms. So autoencoder reconstruction provided a suitable minimalistic setting, free from these influences, to study form-induced representational quantisation., except for an activation function which can selectively reintroduce anisotropy through changes to its form. Hence, this isolates the causality of the phenomenon to the activation functions.
@@
This requires two activation functions, which are highly similar in action but differ in the presence or absence of anisotropy. The isotropic paper Bird (2025a) provides these preliminary isotropic activation functions, which enable this comparative analysis. An analogue of standard elementwise (anisotropic) tanh, displayed in Eqn. 3, can be compared with its orthogonal analogue Isotropic-Tanh, displayed in Eqn. 4 — each is displayed in multivariate form for comparison $\mathbf{f}:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. These functions produce equivalent non-linear maps along the standard basis directions: $f_{\text{iso}}\left(\alpha\hat{e}_{i}\right)=f_{\text{aniso}}\left(\alpha\hat{e}_{i}\right)$ forall $\hat{e}_{i}$ and $\alpha$. Standard-Tanh is algebraically equivariant to the hyperoctahedral group (standard-basis permutation with sign-flips), $B_{n}$, whilst isotropic-tanh is algebraically equivariant to the orthogonal group, $\mathrm{O}\left(n\right)$, with $B_{n}\subset\mathrm{O}\left(n\right)$.
@@
This approach is also later extended to a Leaky-ReLU Maas et al. (2013) family, which are modified to adhere to the following symmetries $S_{n}\subset B_{n}\subset\mathrm{O}\left(n\right)$. This is discussed further in App. B.4 and indicates that algebraic symmetries are the primary predictor, since other analytical properties of Leaky-ReLU substantially differ from the tanh-like family of Eqns. 3 and 4.
@@
Additionally, the PPP method does not produce any angular-falloff smearing effect related to the cone width of SRM. Therefore, this modified method enables higher-fidelity angular maps of representations, which may be beneficial in many circumstances. The following section provides a detailed overview of this tool.
Using a comparative analysis between standard-Tanh and isotropic-Tanh, and similarly on Leaky-ReLU family of functions, this method can then determine whether there is a tendency for discrete representation clusters to form due to the task-agnostic functional form symmetries.